OpenStudy (loser66):

Let f(x,y) =3/2 \(x^2\leq y \leq 1\) \(0\leq x\leq 1\) a)find \(P(0\leq X\leq 1/2)\) b) find \(P (1/2 \leq Y \leq1\) c)Find \(P(X\geq 1/2,Y\geq 1/2)\) Please, help

OpenStudy (loser66):

I got \[f_X(x)= \dfrac{3}{2}(1-x^2)~~~0\leq x\leq 1\] \[f_Y(y) =\dfrac{3}{2}\sqrt{y}~~~0\leq y\leq 1\] and after testing, I have they are dependent. However, I don't know how to do for a, b, c @Zarkon

OpenStudy (loser66):

a) \[\int_0^{1/2}f_X(x) dx=\int_0^{1/2}(3/2)(1-x^2)dx\] right? the same with b, right?

OpenStudy (loser66):

for c) it is  need confirm a, b only

OpenStudy (kirbykirby):

Yup a) and b) look good :) Although the integral bounds for b) will be 1/2 to 1, but I'm pretty sure you knew that ;)

OpenStudy (loser66):

@kirbykirby but I don't know why we have to take \[\int_{1/2}^{1}\int_{1/2}^{1/\sqrt2}(3/2) dxdy + \int_{1/2}^{1}\int_{1/\sqrt 2}^\sqrt y (3/2) dx dy\] I understand the second integral, but the first one. Is it not that it is a rectangular and we just take width x height as \((1/\sqrt 2)-1/2)(1/2)\) ?? why do we have to have that integral?

OpenStudy (kirbykirby):

this is regarding c) right?

OpenStudy (loser66):

yes

OpenStudy (kirbykirby):

ah well it seems that they have broken up the integral into two regions . Maybe the diagram here is more clear  However, it is a uniformly distributed region, so indeed, you can just take the geometric area .. but you must divide that area by the total area of the 2-D region covered by your joint density function